OpenStudy (jennychan12):

Can someone help? d^2y/dx^2(second derivative with respect to x) of xsiny = 3

OpenStudy (jennychan12):

i found that dy/dx = -tany/x so i got d^2y/dx^2 = -tany/x^2

OpenStudy (jennychan12):

is that right?